A commuter train leaves a station( constant operating speed. a) What is the operating speed of the train? b) Over what distance does it achieve that speed? starts from rest) and has a variable acceleration of ) m/s* for the first ten seconds of its motion. Thereafter, the train has a
Answer #1

a(t) = (6 - t)

a) v = \inta dt

v = \int(6 -t) dt

= 6t - t^2 /2

b) distance = \intv dt

= \int ( 6t - t^2 /2) dt

= 3t^2 - t^3 / 6